# On the Brun–Titchmarsh theorem

### v1

@@ 7.2. Initial reductions
$2^{\xi}\gg(\log(MNq))^{A},\ \ \ (\log N)^{\xi}\ll\exp(\xi^{2})\ll(MNq)^{\varepsilon}$
for any $A,\varepsilon>0.$
We first consider a sub-sum
$\mathcal{R}_{0}:=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M,~{}n_{1},n_{2}\in\mathbf{Z}\\ m_{1}n_{1}\equiv m_{2}n_{2}\,(\operatorname{mod}{q})\\ (m_{1}m_{2}n_{1}n_{2},q)=1,~{}\omega(m_{1})>\xi\end{subarray}}\alpha_{m_{1}}\beta_{m_{2}}F(n_{1},n_{2}).$
Trivially, we have
$\mathcal{R}_{0}$ $\ll\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M,~{}n_{1},n_{2}\sim N\\ m_{1}n_{1}\equiv m_{2}n_{2}\,(\operatorname{mod}{q})\\ \omega(m_{1})>\xi\end{subarray}}1\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}m_{1}\sim M,~{}n_{1}\sim N\\ (m_{1}n_{1},q)=1,~{}\omega(m_{1})>\xi\end{subarray}}\sum_{\begin{subarray}{c}MN<k\leqslant 4MN\\ k\equiv m_{1}n_{1}\,(\operatorname{mod}{q})\end{subarray}}\tau(k).$
By Shiu’s bound (see [Sh80, Theorem 2]) we obtain
$\sum_{k}\ll\frac{MN\log(MN)}{\varphi(q)}.$
Moreover,
$\sum_{\begin{subarray}{c}m_{1}\sim M\\ \omega(m_{1})>\xi\end{subarray}}1$ $\leqslant\frac{1}{2^{\xi}}\sum_{m_{1}\sim M}2^{\omega(m)}\ll\frac{M\log M}{2^{\xi}}\ll M(\log MNq)^{-A-2}$
for any $A>0$. Hence
$\mathcal{R}_{0}\ll q^{-1}(MN)^{2}(\log MNq)^{-A}.$
In this way, we may write
(7.5) =~$\mathcal{R}=\widetilde{\mathcal{R}}+O(q^{-1}(MN)^{2}(\log MNq)^{-A}),$
where $\widetilde{\mathcal{R}}$ is defined as $\mathcal{R},$ with an additional restriction that
$\omega(m_{1})>\xi\text{ or }\omega(m_{2})>\xi\Longrightarrow\alpha_{m_{1}}\beta_{m_{2}}=0,$
i.e., we additionally assume that $(m_{1},m_{2})\in\mathscr{F}_{\xi}(2M,2M).$
We now turn to estimate $\widetilde{\mathcal{R}}.$ We first write
~$\widetilde{\mathcal{R}}=\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d,~{}n_{1},n_{2}\in\mathbf{Z}\\ m_{1}n_{1}\equiv m_{2}n_{2}\,(\operatorname{mod}{q})\\ (m_{1}m_{2}n_{1}n_{2},q)=(m_{1},m_{2})=1\end{subarray}}\alpha_{dm_{1}}\beta_{dm_{2}}F(n_{1},n_{2}),$
whence we assume that $(dm_{1},dm_{2})\in\mathscr{F}_{\xi}(2M,2M).$ Note that the congruence restriction $m_{1}n_{1}\equiv m_{2}n_{2}\,(\operatorname{mod}{q})$ is equivalent to
@@
with $R:=2MN/q.$ Therefore,
$\widetilde{\mathcal{R}}$ $=\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\sum_{0\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d\\ (m_{1}m_{2},q)=(m_{1},m_{2})=1\end{subarray}}\alpha_{dm_{1}}\beta_{dm_{2}}$
×$\ \ \ \ \times\sum_{\begin{subarray}{c}n_{1}\equiv qr\overline{m_{1}}\,(\operatorname{mod}{m_{2}})\\ (n_{1},q)=1\end{subarray}}F\Big{(}n_{1},\frac{m_{1}n_{1}-qr}{m_{2}}\Big{)}.$
From Poisson summation, the last $n_{1}$-sum can be transformed as
$\sum_{n_{1}}$ $=\sum_{\delta\mid q}\mu(\delta)\sum_{n_{1}\equiv qr\overline{\delta m_{1}}\,(\operatorname{mod}{m_{2}})}F\Big{(}\delta n_{1},\frac{\delta m_{1}n_{1}-qr}{m_{2}}\Big{)}$
$=\frac{1}{m_{2}}\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{h\in\mathbf{Z}}\widetilde{F}\Big{(}\frac{h}{m_{2}\delta};m_{1},m_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta m_{1}}}{m_{2}}\Big{)},$
@@
$\widetilde{F}(\lambda;m_{1},m_{2},r)=\int_{\mathbf{R}}F\Big{(}y,\frac{m_{1}y-qr}{m_{2}}\Big{)}\mathrm{e}(-\lambda y)\mathrm{d}y.$
@@
(7.6) $\sup_{m_{1},m_{2},r}|\widetilde{F}(\lambda;m_{1},m_{2},r)|$ $\ll N(1+\eta^{-1}|\lambda|N)^{-A}$
@@
$\widetilde{\mathcal{R}}$ $=\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\sum_{0\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d\\ (m_{1}m_{2},q)=(m_{1},m_{2})=1\end{subarray}}\frac{\alpha_{dm_{1}}\beta_{dm_{2}}}{m_{2}}\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{h\in\mathbf{Z}}\widetilde{F}\Big{(}\frac{h}{m_{2}\delta};m_{1},m_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta m_{1}}}{m_{2}}\Big{)}.$
@@
$=\frac{\varphi(q)}{q}\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d\\ (m_{1}m_{2},q)=(m_{1},m_{2})=1\end{subarray}}\frac{\alpha_{dm_{1}}\beta_{dm_{2}}}{m_{2}}\sum_{0\leqslant|r|\leqslant R/d}\widetilde{F}(0;m_{1},m_{2},r).$
@@
$=m_{2}\sum_{r\in\mathbf{Z}}\int_{\mathbf{R}}F\Big{(}m_{2}y,m_{1}y-\frac{qr}{m_{2}}\Big{)}\mathrm{d}y$
$=m_{2}\iint_{\mathbf{R}^{2}}F\Big{(}m_{2}y,m_{1}y-\frac{qz}{m_{2}}\Big{)}\mathrm{d}y\mathrm{d}z+O(\eta q^{-100})$
$=\frac{m_{2}}{q}\iint_{\mathbf{R}^{2}}F(y,z)\mathrm{d}y\mathrm{d}z+O(\eta q^{-100}).$
@@
(7.7) ~$\widetilde{\mathcal{R}}=\mathcal{R}^{*}+\frac{\varphi(q)}{q^{2}}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M\\ (m_{1}m_{2},q)=1\end{subarray}}\alpha_{m_{1}}\beta_{m_{2}}\iint_{\mathbf{R}^{2}}F(y,z)\mathrm{d}y\mathrm{d}z+O(\eta),$
@@
$\mathcal{R}^{*}$ $=\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\sum_{0\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d\\ (m_{1}m_{2},q)=(m_{1},m_{2})=1\end{subarray}}\frac{\alpha_{dm_{1}}\beta_{dm_{2}}}{m_{2}}$
×$\ \ \ \ \times\sum_{h\neq 0}\widetilde{F}\Big{(}\frac{h}{m_{2}\delta};m_{1},m_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta m_{1}}}{m_{2}}\Big{)}.$
@@
$\sum_{h\neq 0}\widetilde{F}\Big{(}\frac{h}{m_{2}\delta};m_{1},m_{2},0\Big{)}$ $\ll\eta N^{-1+\varepsilon}\delta m_{2}$
@@
(7.8) $\mathcal{R}^{*}$ $=\mathcal{R}^{*}_{2}+O(\eta(MN)^{1+\varepsilon}),$
@@
$\mathcal{R}^{*}_{2}$ $=\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\sum_{1\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d\\ (m_{1}m_{2},q)=(m_{1},m_{2})=1\end{subarray}}\frac{\alpha_{dm_{1}}\beta_{dm_{2}}}{m_{2}}$
×$\ \ \ \ \times\sum_{1\leqslant|h|\leqslant H\delta/d}\widetilde{F}\Big{(}\frac{h}{m_{2}\delta};m_{1},m_{2},r\Big{)}\mathrm{e}\Big{(}\frac{(q/\delta)hr\overline{m_{1}}}{m_{2}}\Big{)}$
with $H:=\eta M^{1+\varepsilon}/N.$
@@ 7.3. Estimates for exponential sums in $\mathcal{R}^{*}_{2}$
$\mathcal{R}^{*}_{2}$ $\ll(MNq)^{\varepsilon}\sum_{\delta\mid q}\frac{1}{\delta}\sum_{\begin{subarray}{c}d_{1}d_{2}=d\leqslant 2M\\ (d,q)=1\end{subarray}}\sup|\mathcal{R}^{*}_{2}(A_{1},B_{1},A_{2},B_{2},R^{\prime};d_{1},d_{2},\delta)|,$
@@
$1\leqslant A_{1}\leqslant\frac{M_{1}}{d_{1}},~{}1\leqslant B_{1}\leqslant\frac{N_{1}}{d_{2}},~{}1\leqslant A_{2}\leqslant\frac{M_{2}}{d_{1}},~{}1\leqslant B_{2}\leqslant\frac{N_{2}}{d_{2}},~{}1\leqslant R^{\prime}\leqslant\frac{R}{d},$
@@
$\mathcal{R}^{*}_{2}(\cdots)$ $=\sum_{|r|\sim R^{\prime}}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1}\sim A_{1},~{}n_{1}\sim B_{1},~{}m_{2}\sim A_{2},~{}n_{2}\sim B_{2}\\ (m_{1}m_{2}n_{1}n_{2},q)=(m_{1}n_{1},m_{2}n_{2})=1\end{subarray}}\frac{\alpha^{\prime}_{d_{1}m_{1}}\alpha^{\prime\prime}_{d_{2}n_{1}}\beta^{\prime}_{d_{1}m_{2}}\beta^{\prime\prime}_{d_{2}n_{2}}}{m_{2}n_{2}}$
×$\ \ \ \ \times\sum_{1\leqslant|h|\leqslant H\delta/d}\widetilde{F}\Big{(}\frac{h}{m_{2}n_{2}\delta};m_{1}n_{1},m_{2}n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{(q/\delta)hr\overline{m_{1}n_{1}}}{m_{2}n_{2}}\Big{)}.$
@@
$=m_{2}n_{2}\delta\int_{\frac{dN}{2M\delta}}^{\frac{2dN}{M\delta}}F\Big{(}m_{2}n_{2}y\delta,m_{1}n_{1}y\delta-\frac{qr}{m_{2}n_{2}}\Big{)}\mathrm{e}(-hy)\mathrm{d}y$
$=m_{2}n_{2}\delta\int_{\frac{dN}{2M\delta}}^{\frac{2dN}{M\delta}}\mathrm{e}(-hy)\mathrm{d}y\iint_{\mathbf{R}^{2}}\widehat{F}(y_{1},y_{2})\mathrm{e}\Big{(}m_{2}n_{2}y_{1}y\delta+m_{1}n_{1}y_{2}y\delta-\frac{qry_{2}}{m_{2}n_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}$
$=\frac{m_{2}n_{2}\delta}{r^{2}}\int_{\frac{dN}{2M\delta}}^{\frac{2dN}{M\delta}}\mathrm{e}(-hy)\mathrm{d}y\iint_{\mathbf{R}^{2}}\widehat{F}\Big{(}\frac{m_{1}y_{1}}{n_{2}r},\frac{y_{2}n_{2}}{m_{1}r}\Big{)}\mathrm{e}\Big{(}\frac{(m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta}{r}-\frac{qy_{2}}{m_{1}m_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}$
$=\frac{m_{2}n_{2}\delta}{r}\int_{\frac{dN}{2Mr\delta}}^{\frac{2dN}{Mr\delta}}\mathrm{e}(-hry)\mathrm{d}y\iint_{\mathbf{R}^{2}}\widehat{F}\Big{(}\frac{m_{1}y_{1}}{n_{2}r},\frac{y_{2}n_{2}}{m_{1}r}\Big{)}\mathrm{e}\Big{(}(m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}.$
@@
(7.9) $\frac{\partial^{\nu_{1}+\nu_{2}}}{\partial y_{1}^{\nu_{1}}\partial y_{2}^{\nu_{2}}}\widehat{F}(y_{1},y_{2})\ll N^{\nu_{1}+\nu_{2}+2}(1+\eta^{-1}N|y_{1}|)^{-c_{1}}(1+\eta^{-1}N|y_{2}|)^{-c_{2}}$
@@
$U=\frac{1}{M^{2}N^{4}},\ \ V_{1}=\frac{\eta R^{\prime}B_{2}}{A_{1}N}N^{\varepsilon}.$
@@
$\ll\frac{\eta\delta HA_{1}^{2}A_{2}B_{1}}{M^{3}N^{2}}\ll\frac{\eta^{2}\delta M^{\varepsilon}}{M^{2}N},$
@@
$V_{2}=\frac{\eta R^{\prime}A_{1}}{B_{2}N}N^{\varepsilon}.$
@@
$=\frac{m_{2}n_{2}\delta}{4\pi^{2}r}\int_{\frac{dN}{4MR^{\prime}\delta}}^{\frac{4dN}{MR^{\prime}\delta}}\mathrm{e}(-hry)\mathrm{d}y\int_{U}^{V_{1}}\int_{U}^{V_{2}}\mathrm{e}\Big{(}(m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}$

### v3

@@ 7.2. Initial reductions
$\mathcal{R}=\sum_{\begin{subarray}{c}d\leqslant 2N\\ (d,q)=1\end{subarray}}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/d,~{}l_{1},l_{2}\in\mathbf{Z}\\ n_{1}l_{1}\equiv n_{2}l_{2}\,(\operatorname{mod}{q})\\ (n_{1}n_{2}l_{1}l_{2},q)=(n_{1},n_{2})=1\end{subarray}}\alpha_{dn_{1}}\beta_{dn_{2}}F(l_{1},l_{2}).$
Note that the congruence restriction $n_{1}l_{1}\equiv n_{2}l_{2}\,(\operatorname{mod}{q})$ is equivalent to
@@
with $R:=2NL/q.$ Therefore,
$\mathcal{R}$ $=\sum_{\begin{subarray}{c}d\leqslant 2N\\ (d,q)=1\end{subarray}}\sum_{0\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/d\\ (n_{1}n_{2},q)=(n_{1},n_{2})=1\end{subarray}}\alpha_{dn_{1}}\beta_{dn_{2}}\sum_{\begin{subarray}{c}l_{1}\equiv qr\overline{n_{1}}\,(\operatorname{mod}{n_{2}})\\ (l_{1},q)=1\end{subarray}}F\Big{(}l_{1},\frac{n_{1}l_{1}-qr}{n_{2}}\Big{)}.$
From Poisson summation, the last $l_{1}$-sum can be transformed as
$\sum_{l_{1}}$ $=\sum_{\delta\mid q}\mu(\delta)\sum_{l_{1}\equiv qr\overline{\delta n_{1}}\,(\operatorname{mod}{n_{2}})}F\Big{(}\delta l_{1},\frac{\delta n_{1}l_{1}-qr}{n_{2}}\Big{)}$
$=\frac{1}{n_{2}}\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{h\in\mathbf{Z}}\widetilde{F}\Big{(}\frac{h}{n_{2}\delta};n_{1},n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta n_{1}}}{n_{2}}\Big{)},$
@@
$\widetilde{F}(\lambda;n_{1},n_{2},r)=\int_{\mathbf{R}}F\Big{(}y,\frac{n_{1}y-qr}{n_{2}}\Big{)}\mathrm{e}(-\lambda y)\mathrm{d}y.$
@@
(7.5) $\sup_{n_{1},n_{2},r}|\widetilde{F}(\lambda;n_{1},n_{2},r)|$ $\ll L(1+\eta^{-1}|\lambda|L)^{-A}$
@@
$\mathcal{R}$ $=\sum_{\begin{subarray}{c}d\leqslant 2N\\ (d,q)=1\end{subarray}}\sum_{0\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/d\\ (n_{1}n_{2},q)=(n_{1},n_{2})=1\end{subarray}}\frac{\alpha_{dn_{1}}\beta_{dn_{2}}}{n_{2}}\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}$
×$\ \ \ \ \times\sum_{h\in\mathbf{Z}}\widetilde{F}\Big{(}\frac{h}{n_{2}\delta};n_{1},n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta n_{1}}}{n_{2}}\Big{)}.$
@@
(7.6) $=\frac{\varphi(q)}{q}\sum_{\begin{subarray}{c}d\leqslant 2N\\ (d,q)=1\end{subarray}}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/d\\ (n_{1}n_{2},q)=(n_{1},n_{2})=1\end{subarray}}\frac{\alpha_{dn_{1}}\beta_{dn_{2}}}{n_{2}}\sum_{0\leqslant|r|\leqslant R/d}\widetilde{F}(0;n_{1},n_{2},r).$
@@
$=n_{2}\sum_{r\in\mathbf{Z}}\int_{\mathbf{R}}F\Big{(}n_{2}y,n_{1}y-\frac{qr}{n_{2}}\Big{)}\mathrm{d}y=n_{2}\iint_{\mathbf{R}^{2}}F\Big{(}n_{2}y,n_{1}y-\frac{qz}{n_{2}}\Big{)}\mathrm{d}y\mathrm{d}[z].$
The above Stieltjes integral is equal to
$=n_{2}\iint_{\mathbf{R}^{2}}F\Big{(}n_{2}y,n_{1}y-\frac{qz}{n_{2}}\Big{)}\mathrm{d}y\mathrm{d}z-q\iint_{\mathbf{R}^{2}}\{z\}F^{\prime}\Big{(}n_{2}y,n_{1}y-\frac{qz}{n_{2}}\Big{)}\mathrm{d}y\mathrm{d}z,$
where $F^{\prime}(u,v)=\frac{\partial}{\partial v}F(u,v)$. The second term contributes to (7.6) at most $O(\eta(NL)^{1+\varepsilon}),$ and the first term is exactly
$\frac{n_{2}}{q}\iint_{\mathbf{R}^{2}}F(y,z)\mathrm{d}y\mathrm{d}z.$
@@
(7.7) $\mathcal{R}=\mathcal{R}^{*}+\frac{\varphi(q)}{q^{2}}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N\\ (n_{1}n_{2},q)=1\end{subarray}}\alpha_{n_{1}}\beta_{n_{2}}\iint_{\mathbf{R}^{2}}F(y,z)\mathrm{d}y\mathrm{d}z+O(\eta(NL)^{1+\varepsilon}),$
@@
$\mathcal{R}^{*}$ $=\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{\begin{subarray}{c}d\leqslant 2N\\ (d,q)=1\end{subarray}}\sum_{0\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/d\\ (n_{1}n_{2},q)=(n_{1},n_{2})=1\end{subarray}}\frac{\alpha_{dn_{1}}\beta_{dn_{2}}}{n_{2}}$
×$\ \ \ \ \times\sum_{h\neq 0}\widetilde{F}\Big{(}\frac{h}{n_{2}\delta};n_{1},n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta n_{1}}}{n_{2}}\Big{)}.$
@@
$\sum_{h\neq 0}\widetilde{F}\Big{(}\frac{h}{n_{2}\delta};n_{1},n_{2},0\Big{)}$ $\ll\eta L^{-1+\varepsilon}\delta n_{2}$
@@
(7.8) $\mathcal{R}^{*}$ $=\mathcal{R}^{*}_{2}+O(\eta(NL)^{1+\varepsilon}),$
@@
(7.9) $\begin{split}\mathcal{R}^{*}_{2}&=\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{\begin{subarray}{c}d\leqslant 2N\\ (d,q)=1\end{subarray}}\sum_{1\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/d\\ (n_{1}n_{2},q)=(n_{1},n_{2})=1\end{subarray}}\frac{\alpha_{dn_{1}}\beta_{dn_{2}}}{n_{2}}\\ &\ \ \ \ \times\sum_{h\neq 0}\widetilde{F}\Big{(}\frac{h}{n_{2}\delta};n_{1},n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta n_{1}}}{n_{2}}\Big{)}\end{split}$
@@ 7.3. Estimates for exponential sums in $\mathcal{R}^{*}_{2}$
$\mathcal{R}^{*}_{2}$ $\ll(qNL)^{\varepsilon}\sum_{\delta\mid q}\frac{1}{\delta}\sum_{\begin{subarray}{c}d_{1}d_{2}=d_{1}^{\prime}d_{2}^{\prime}=d\leqslant 2N\\ (d,q)=1\end{subarray}}\sup|\mathcal{R}^{*}_{2}(A_{1},B_{1},A_{2},B_{2},R^{\prime};d_{1},d_{2},d_{1}^{\prime},d_{2}^{\prime},\delta)|+1,$
@@
$1\leqslant A_{1}\leqslant\frac{M_{1}}{d_{1}},~{}1\leqslant B_{1}\leqslant\frac{N_{1}}{d_{2}},~{}1\leqslant A_{2}\leqslant\frac{M_{2}}{d_{1}^{\prime}},~{}1\leqslant B_{2}\leqslant\frac{N_{2}}{d_{2}^{\prime}},~{}1\leqslant R^{\prime}\leqslant\frac{R}{d},$
@@
$\mathcal{R}^{*}_{2}(\cdots)$ $=\sum_{|r|\sim R^{\prime}}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1}\sim A_{1},~{}n_{1}\sim B_{1},~{}m_{2}\sim A_{2},~{}n_{2}\sim B_{2}\\ (m_{1}m_{2}n_{1}n_{2},q)=(m_{1}n_{1},m_{2}n_{2})=(m_{1},d_{2})=(m_{2},d_{2}^{\prime})=1\end{subarray}}\frac{\alpha^{\prime}_{d_{1}m_{1}}\alpha^{\prime\prime}_{d_{2}n_{1}}\beta^{\prime}_{d_{1}^{\prime}m_{2}}\beta^{\prime\prime}_{d_{2}^{\prime}n_{2}}}{m_{2}n_{2}}$
×$\ \ \ \ \times\sum_{1\leqslant|h|\leqslant H}\widetilde{F}\Big{(}\frac{h}{m_{2}n_{2}\delta};m_{1}n_{1},m_{2}n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{(q/\delta)hr\overline{m_{1}n_{1}}}{m_{2}n_{2}}\Big{)}.$
Here we take $H=\eta\delta A_{2}B_{2}L^{-1+\varepsilon}.$
@@
$=m_{2}n_{2}\delta\int_{\frac{rL}{8A_{2}B_{2}R^{\prime}\delta}}^{\frac{2rL}{A_{2}B_{2}R^{\prime}\delta}}F\Big{(}m_{2}n_{2}y\delta,m_{1}n_{1}y\delta-\frac{qr}{m_{2}n_{2}}\Big{)}\mathrm{e}(-hy)\mathrm{d}y$
$=m_{2}n_{2}\delta\int_{\frac{rL}{8A_{2}B_{2}R^{\prime}\delta}}^{\frac{2rL}{A_{2}B_{2}R^{\prime}\delta}}\mathrm{e}(-hy)\mathrm{d}y\iint_{\mathbf{R}^{2}}\widehat{F}(y_{1},y_{2})\mathrm{e}\Big{(}m_{2}n_{2}y_{1}y\delta+m_{1}n_{1}y_{2}y\delta-\frac{qry_{2}}{m_{2}n_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}$
$=\frac{m_{2}n_{2}\delta}{r^{2}}\int_{\frac{rL}{8A_{2}B_{2}R^{\prime}\delta}}^{\frac{2rL}{A_{2}B_{2}R^{\prime}\delta}}\mathrm{e}(-hy)\mathrm{d}y\iint_{\mathbf{R}^{2}}\widehat{F}\Big{(}\frac{m_{1}y_{1}}{n_{2}r},\frac{y_{2}n_{2}}{m_{1}r}\Big{)}\mathrm{e}\Big{(}\frac{(m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta}{r}-\frac{qy_{2}}{m_{1}m_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}$
$=\frac{m_{2}n_{2}\delta}{r}\int_{\frac{L}{8A_{2}B_{2}R^{\prime}\delta}}^{\frac{2L}{A_{2}B_{2}R^{\prime}\delta}}\mathrm{e}(-hry)\mathrm{d}y\iint_{\mathbf{R}^{2}}\widehat{F}\Big{(}\frac{m_{1}y_{1}}{n_{2}r},\frac{y_{2}n_{2}}{m_{1}r}\Big{)}\mathrm{e}\Big{(}(m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}.$
@@
(7.10) $\frac{\partial^{\nu_{1}+\nu_{2}}}{\partial y_{1}^{\nu_{1}}\partial y_{2}^{\nu_{2}}}\widehat{F}(y_{1},y_{2})\ll L^{\nu_{1}+\nu_{2}+2}(1+\eta^{-1}L|y_{1}|)^{-c_{1}}(1+\eta^{-1}L|y_{2}|)^{-c_{2}}$
@@
$U=\frac{1}{N^{3}L},\ \ V_{1}=\frac{\eta R^{\prime}B_{2}}{A_{1}L}L^{\varepsilon}.$
@@
$\ll\frac{\eta^{2}\delta N^{\varepsilon}}{N^{3}}\frac{A_{1}^{2}B_{1}}{B_{2}}\ll\eta^{2}\delta d^{-1}N^{\varepsilon},$
@@
$V_{2}=\frac{\eta R^{\prime}A_{1}}{B_{2}L}L^{\varepsilon}.$
@@
$=\frac{m_{2}n_{2}\delta}{4\pi^{2}r}\int_{\frac{L}{8A_{2}B_{2}R^{\prime}\delta}}^{\frac{2L}{A_{2}B_{2}R^{\prime}\delta}}\mathrm{e}(-hry)\mathrm{d}y\int_{U}^{V_{1}}\int_{U}^{V_{2}}\mathrm{e}\Big{(}(m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}$